Find the inverse of the function f(x)= 4/x+7, what is the range of f(x)?
Answer by uma(370)

f(x) = 4/(x+7)
that is y = 4/(x+7)
First we interchange x and y
==> x = 4/(y+7)
==> x(y+7) = 4
==> xy + 7x = 4
==> xy = 4 - 7x
==> xy/x = (4-7x)/x
==> y = (4-7x)/x
The inverse of the given function is f^(-1) (x) = (4-7x)/x

We see the above is not defined at x = 0
So range of the given function is all real numbers except 0.
Range = R - {0}

Good Luck!!!